Register Login Password forgotten?

Station History Log - Potsdam, Germany (7841)

SOD Year Doy Toy System Description
78418701 2017 062 12:00 06.01 A033-ET Event Timer replacing A032-ET
78418701 2016 362 16:00 04.01 New daylight filter with higher peak transmission inserted
78418701 2016 175 12:00 05.02 Pump diode module for regenerative amplifier exchanged; 10 dB attenuator re-inserted into start channel (change in calibration mean value)
78418701 2016 145 14:00 04.01 Optical components of telescopes cleaned (negative achromat Tx, entrance window and coupling mirror Rx)
78418701 2016 083 12:00 05.02 Repaired pump diode for external amplifier and new coupling fibre installed, maximum output power 530 mW
78418701 2016 056 17:00 05.02 Pump diode for external amplifier and coupling fibre failed, laser operations continued with minimum output power (50 mW)
78418701 2016 033 10:00 05.02 Repaired HV power supply for Pockels cell inserted, operations resumed
78418701 2016 018 06:00 05.02 HV power supply for Pockels cell failed, tracking suspended
78418701 2015 296 10:00 04.01 Interpolator unit in Rx azimuth axis replaced, tracking resumed
78418701 2015 291 20:00 04.01 Interpolator unit in Rx azimuth axis faulty, tracking suspended
78418701 2015 204 10:00 05.02 Drop in start photodiode amplitude, 10 dB attenuater removed from start channel (change in target mean value)
78418701 2014 246 08:00 05.02 Controller and coupling fibre exchanged, operations resumed with nominal output power
78418701 2014 218 09:00 05.02 Controller of oscillator stage failed, station operations interrupted
78418701 2014 125 14:00 05.02 Coupling fibre for final amplifier stage broken, output reduced to 40 mW average power
78418701 2013 264 16:00 04.01 Azimuth encoder of Tx telescope exchanged
78418701 2013 145 12:00 04.01 Narrowband daylight filter and wideband (3 nm) nighttime filter selectable for signal optimization
78418701 2012 341 16:00 05.02 Mating and de-mating of laser control cables to restore proper TEC operation
78418701 2012 270 16:00 06.01 Faulty PMT H5320 replaced by spare unit of same type
78418701 2012 089 12:00 05.02 Experimental start pulse converter removed
78418701 2012 088 12:00 05.02 Start pulse converter with constant fraction properties inserted for test
78418701 2012 039 09:00 06.01 PMT operational voltage raised from 1600 to 1700 V
78418701 2012 011 12:00 05.02 Coupling mirror inside the laser replaced by 25-mm HR type
78418701 2011 314 12:00 05.02 Too high sensitivity of start pulse converter detected, 10 dB attenuator inserted
78418701 2011 242 12:00 05.02 kHz Nd:YVO laser replacing 10 Hz Nd:YAG system
78418701 2011 242 12:00 06.01 A032-ET Event Timer replacing A031-ET
78418701 2011 242 12:00 06.02 MPD-1CTC SPAD replacing AD230 SPAD
78418701 2008 300 00:00 04.01 Daylight filter replaced by same type and FWHM
78418701 2004 050 00:00 06.01 A031 Event timer replacing SR620 time interval counter
78418701 2003 150 00:00 04.01 Daylight filter (0.4 nm FWHM) introduced